9. Find the symmetric matrix A associated with the given quadratic form. A. x +223 +63122B. 2x2 – 3y2 + x2 - 4x2
Answer #1

9 А. x² + 2x² - 6xx X X₂ A= 1 3 [X;X27 3 2 3 2 o B : 23357+22_axz х 2 9 Z Be х -2 ។ -3 O Z -2 . = [Xyz 2 2 0-3 fan -2 So A =